# Grasping a Handful: Sequential Multi-Object Dexterous Grasp Generation

### v3

@@ Abstract
We introduce the sequential multi-object robotic grasp sampling algorithm SeqGrasp that can robustly synthesize stable grasps on diverse objects using the robotic hand’s partial Degrees of Freedom (DoF). We use SeqGrasp to construct the large-scale Allegro Hand sequential grasping dataset SeqDataset and use it for training the diffusion-based sequential grasp generator SeqDiffuser. We experimentally evaluate SeqGrasp and SeqDiffuser against the state-of-the-art non-sequential multi-object grasp generation method MultiGrasp in simulation and on a real robot. The experimental results demonstrate that SeqGrasp and SeqDiffuser reach an 8.71%-43.33% higher grasp success rate than MultiGrasp. Furthermore, SeqDiffuser is approximately 1000 times faster at generating grasps than SeqGrasp and MultiGrasp.
@@ I Introduction
Dexterous grasping has been extensively studied in robotics [1, 2]. The field has evolved through two primary research directions. First, numerous methods have been developed for generating robotic grasps [3, 4, 5, 6, 7, 8], from analytical approaches [7] to data-driven generative models including Generative Adversarial Networks [5], diffusion models [4, 3], and Variational Autoencoders [8, 6]. Second, a complementary line of research has focused on understanding human grasping behaviors, including intention recognition [9], real-time tracking of hand-object interactions [10], grasp taxonomy classification [11], and learning task-oriented grasping strategies from human demonstrations [12]. The primary strategy in dexterous grasp generation has been to bring the robotic hand close to the object and then simultaneously envelop it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [7].
@@ II Related Work
MOG [14] Human&Barrett Sim./Real 11 \ 28K Simu./Seq. \
Grasp’Em [13] Shadow Sim. 8 36 90K Simultaneous 2
SeqDataset(Ours) Allegro Sim. 509 2400 870K Sequential 4 ✓*
TABLE I: Multi-object grasping datasets. *Our dataset will be made publicly available upon acceptance.
@@ II-C Datasets for Dexterous Grasping
Large-scale dexterous grasp datasets [5, 7, 26, 14, 13, 21, 27, 28] have significantly advanced the training of data-driven dexterous grasping methods. However, most of these datasets target single-object grasping [5, 7, 26, 21, 27, 28], with only a few for multi-object grasping [14, 13]. As shown in Table I, these existing multi-object datasets are small and predominantly focus on simultaneous rather than sequential grasping. Therefore, we collect the new large-scale sequential multi-object grasping dataset SeqDataset using our method SeqGrasp. SeqDataset is, to date, by far the largest multi-object grasping dataset.
@@ IV-C Energy Function
where $\mathbf{I}$ represents the identity matrix, and $[\mathbf{x}_{j}]_{\times}$ (${1\leq j\leq 2}$) denotes the skew-symmetric matrix formed from the contact point $\mathbf{x}_{j}$.
The contact distance and penetration terms ($E_{\text{dis}}$ $E_{\text{hop}})$ encourage the hand-object contacts to occur close to the object surface without penetrating it. The contact distance is mathematically defined as
@@
The self-collision and joint limit terms ($E_{\text{hsp}}$ $E_{\text{joint}}$) encourage physical feasibility. We define these as
@@ V Dataset Generation
We use SeqGrasp to generate our large-scale dataset SeqDataset containing 4.9 million sequential Allegro Hand grasps on over 509 objects from DexGraspNet [7]. Each object is resized to fit within the Allegro Hand by scaling its axis-aligned bounding box to be between 0.06 and 0.10 meters. Then, from the 509 resized objects, we generate 600 unique object sets containing four objects each. The objects in these sets are randomly permuted four times, resulting in 2,400 unique object sequences.
[…]
@@ VI Conditional Sequential Grasp Diffuser
$q(\mathbf{\tilde{g}}{{}_{t}}|\mathbf{\tilde{g}}{{}_{t-1}})$ $=\mathcal{N}(\mathbf{\tilde{g}}{{}_{t}};\mathbf{k}\odot\sqrt{1-\beta_{t}}\mathbf{\tilde{g}}{{}_{t-1}},\beta_{t}\mathbf{I}),$ (12)
@@ VII-A Simulation Experiments
TABLE III: Simulation results. MulG, SeqG, SeqD, G’Em, and SData are short for the MultiGrasp, SeqGrasp, SeqDiffuser, Grasp’Em, and SeqDataset, respectively. The $-i$ following the method name denotes the number of objects used for grasp generation. $\uparrow$($\downarrow$), the higher (lower), the better.
@@ VII-B Real-World Experiments
max width= Real exp. SeqG-3 SeqG-4 SeqD-3 SeqD-4 MulG-3 MulG-4 Printed Obj. Succ. trials 5/10 5/10 4/10 0/10 1/10 1/7 Succ. objects 2.3$\pm$0.9/3 3.4$\pm$0.8/4 2.1$\pm$0.9/3 1.5$\pm$0.8/4 1.2$\pm$1.1/3 1.6$\pm$1.5/4 Real Obj. Succ. trials 5/10 / 0/10 / 0/6 / Succ. objects 2.1$\pm$1.2/3 / 1.8$\pm$0.4/3 / 1.0$\pm$0.6/3 /
TABLE IV: Real-world experimental results. Each entry reports the number of successful trials out of 10 and the average number of objects successfully grasped in mean±$\pm$±std after the test.
The experimental results are presented in Table IV. Similar to the simulation results, SeqGrasp outperformed SeqDiffuser and MultiGrasp in both three-object and four-object grasping tasks, achieving an average success rate of 50%. The low success rate of SeqDiffuser (13%) is primarily due to substantial object inter-penetration in the generated grasps, which meant that the closest physically feasible grasp replicable on the real hand differed substantially from the intended configuration. MultiGrasp performed worst, with an average success rate of 8.7%. The reason is its reliance on object-object contacts for grasp stability, which are sensitive to minor object displacements. This resulted in frequent failures, particularly in four-object grasping trials where MultiGrasp achieved no successful trials and failed to generate valid grasps in three of ten attempts (last column in Table IV). Accounting for these additional failed trials, the average success rate for MultiGrasp decreases to 6.7%.

### v4

@@ Abstract
We introduce the sequential multi-object robotic grasp sampling algorithm SeqGrasp that can robustly synthesize stable grasps on diverse objects using the robotic hand’s partial Degrees of Freedom (DoF). We use SeqGrasp to construct the large-scale Allegro Hand sequential grasping dataset SeqDataset and use it for training the diffusion-based sequential grasp generator SeqDiffuser. We experimentally evaluate SeqGrasp and SeqDiffuser against the state-of-the-art non-sequential multi-object grasp generation method MultiGrasp in simulation and on a real robot. The experimental results demonstrate that SeqGrasp and SeqDiffuser reach an 8.71%-43.33% higher grasp success rate than MultiGrasp. Furthermore, SeqDiffuser is approximately 1000 times faster at generating grasps than SeqGrasp and MultiGrasp. Project page: https://yulihn.github.io/SeqGrasp/.
@@ I Introduction
Dexterous grasping has been extensively studied in robotics [1, 2]. The field has evolved through two primary research directions. First, numerous methods have been developed for generating robotic grasps [3, 4, 5, 6, 7, 8], from analytical approaches [7] to data-driven generative models including Generative Adversarial Networks [5], diffusion models [4, 3], and Variational Autoencoders [8, 6]. Second, a complementary line of research has focused on understanding human grasping behaviors, including intention recognition [9], real-time tracking of hand-object interactions [10], grasp taxonomy classification [11], and learning task-oriented grasping strategies from human demonstrations [12].
The primary strategy in dexterous grasp generation has been to bring the robotic hand close to the object and then simultaneously envelop it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [7].
@@
MOG [14] Human&Barrett Sim./Real 11 \ 28K Simu./Seq. \
Grasp’Em [13] Shadow Sim. 8 36 90K Simultaneous 2
SeqDataset(Ours) Allegro Sim. 509 2400 870K Sequential 4
TABLE I: Multi-object grasping datasets.
@@ II-C Datasets for Dexterous Grasping
Large-scale dexterous grasp datasets [5, 7, 26, 14, 13, 21, 27, 28] have significantly advanced the training of data-driven dexterous grasping methods. However, most of these datasets target single-object grasping [5, 7, 26, 21, 27, 28], with only a few for multi-object grasping [14, 13]. As shown in Table I, these existing multi-object datasets are small and predominantly focus on simultaneous rather than sequential grasping. Therefore, we collect the new large-scale sequential multi-object grasping dataset SeqDataset [^1] The dataset is available at https://yulihn.github.io/SeqGrasp/. using our method SeqGrasp. SeqDataset is, to date, by far the largest multi-object grasping dataset.
@@ IV-C Energy Function
where $\mathbf{I}$ represents the identity matrix, and $\mathbf{x}_{j}]_{\times}$ ($1\leq j\leq 2$}) denotes the skew-symmetric matrix formed from the contact point $\mathbf{x}_{j}$.
The contact distance and penetration terms ($E_{\text{dis}}$ & $E_{\text{hop}})$ encourage the hand-object contacts to occur close to the object surface without penetrating it. The contact distance is mathematically defined as
@@
The self-collision and joint limit terms ($E_{\text{hsp}}$ & $E_{\text{joint}}$) encourage physical feasibility. We define these as
@@ V Dataset Generation
We use SeqGrasp to generate our large-scale dataset SeqDataset  [^2] The dataset is available at https://yulihn.github.io/SeqGrasp/. containing 4.9 million sequential Allegro Hand grasps on over 509 objects from DexGraspNet [7]. Each object is resized to fit within the Allegro Hand by scaling its axis-aligned bounding box to be between 0.06 and 0.10 meters. Then, from the 509 resized objects, we generate 600 unique object sets containing four objects each. The objects in these sets are randomly permuted four times, resulting in 2,400 unique object sequences.
We run Algorithm 1 on the 2,400 unique object sequences with $\mathbf{w}=[50,50,w_{\text{hop}},5,1,5]^{T}$, where $w_{\text{hop}}$, that penalizes hand-object penetration, grows linearly from $5$ to $5e^{2}$. The optimization runs for 6,000 iterations per grasp. We validate the generated grasp sequences in the physics simulator Isaac Gym following the setup in [7]. This setup initializes all object densities to 500 $\text{kg/m}^{3}$, the friction coefficient to 2.0, the hand to the generated grasp configuration, and objects as free-floating. Then, the hand closes until contact is established with the object, which happens when the distance between the hand and the object is less than 2 mm. Finally, a grasp stability test is evaluated by linearly applying an acceleration of 9.8 $\text{m}/\text{s}^{2}$ in all six orthogonal directions for 100 consecutive simulation steps. A grasp sequence is successful if all grasped objects remain in contact with the hand after the grasp stability test and the maximum penetration depth is less than 1 cm. Otherwise, the grasp sequence is a failure. This procedure ultimately produced 4.9 million grasps, of which 870K (17.82%) were successful. Only successful grasps were retained in SeqDataset. The entire data generation process requires approximately 2,500 GPU hours on an NVIDIA A100.
@@ VI Conditional Sequential Grasp Diffuser
[…]
@@ VII-A Simulation Experiments
TABLE III: Simulation results. MulG, SeqG, SeqD, G’Em, and SData are short for the MultiGrasp, SeqGrasp, SeqDiffuser, Grasp’Em, and SeqDataset, respectively. The i-i following the method name denotes the number of objects used for grasp generation. $\uparrow$($\downarrow$), the higher (lower), the better.
@@ VII-B Real-World Experiments
[…]
The experimental results are presented in Table IV. Similar to the simulation results, SeqGrasp outperformed SeqDiffuser and MultiGrasp in both three-object and four-object grasping tasks, achieving an average success rate of 50%. The low success rate of SeqDiffuser (13%) is primarily due to substantial object inter-penetration in the generated grasps, which meant that the closest physically feasible grasp replicable on the real hand differed substantially from the intended configuration. MultiGrasp performed worst, with an average success rate of 8.7%. The reason is its reliance on object-object contacts for grasp stability, which are sensitive to minor object displacements. This resulted in frequent failures, particularly in four-object grasping trials where MultiGrasp achieved no successful trials and failed to generate valid grasps in three of ten attempts (last column in Table IV). Accounting for these additional failed trials lowers the average success rate for MultiGrasp from 8.7% to 6.7%.